OpenStudy (anonymous):

What is the solution to x/4 + 5 = 6? A. 44 B. 2.75 C. 4 D. 0.25

OpenStudy (nurali):

x/4 + 5 = 6 x/4 =6-5 x/4 =1 x=1*4

OpenStudy (anonymous):

so th3 answ3r is C?

OpenStudy (nurali):

Yes